16. A store sells ice cream. Each ice cream contains two tablespoons of ice cream. Possible flavors are vanilla, strawberry, chocolate and mint. Both spoons can be the same taste or different flavors. In one hour the store sells 180 tablespoons of ice cream. The pie chart presents the following data: Chocolate = 100ok, vanilla = 90⁰, strawberries = 150 lu, mint = x⁰. How many ice cream scoops of each flavor sold are thrown away in an hour?


1
Expert's answer
2021-06-01T11:49:23-0400
"100\degree+90\degree+150\degree+x\degree =360\degree"

"x\degree=20\degree"

"chocolate =\dfrac{100\degree}{360\degree}\cdot180=50"

"vanilla =\dfrac{90\degree}{360\degree}\cdot180=45"

"strawberries =\dfrac{150\degree}{360\degree}\cdot180=75"

"mint =\dfrac{20\degree}{360\degree}\cdot180=10"